# Fusion Systems related to polynomial representations of $\operatorname{SL}_{2}(q)$

### v1

@@ 2. The structure of polynomial $p$-groups and their automorphisms
Observe that $C_{G}(C_{V_{n}(q)}(T))\leq N_{G}(T)$ and $|N_{G}(T)|=q.(q-1)$. Clearly, $T$ centralises $C_{V_{n}(q)}(T)$. Let $K$ be a Hall $p^{\prime}$-subgroup of $N_{G}(T)$ so that $K$ is cyclic of order $q-1$ and let $k$ be a generator for $K$. Then we can arrange that $k$ acts as $\lambda^{n}$ on $C_{V_{n}(q)}(T)$ for some $\lambda\in\mathbb{K}$ of multiplicative order $q-1$. Thus every element of $K$ acts as $\mu^{n}$ on $C_{V_{n}(q)}(T)$ for some distinct $\mu\in\mathbb{K}^{*}$. Hence, the elements of $K$ which centralise $C_{V_{n}(q)}(T)$ are in bijection with the $n$th roots of unity in $\mathbb{K}$, which have order $(n,q-1)$ and the claim holds. ∎
For $n\leqslant p-1$ and $0\neq f=\displaystyle\sum_{i=0}^{n}a_{i}x^{n-i}y^{i}\in V$ define $\mathrm{wt}(f)=\mathrm{max}\{i\mid a_{i}\neq 0\}$ to be the weight of $f$. Also define $\mathrm{wt}(0)=-1$. For $-1\leqslant i\leqslant n$, set $C_{i}=\{f\in V\mid\mathrm{wt}(f)\leqslant i\}.$ For each $-1\leqslant i\leqslant n$, $C_{i}$ is a subgroup of $V_{n}(q)$ of order $q^{i+1}$. In particular, $C_{-1}=1$ and $C_{n}=V$.
@@
Suppose that $11\leqslant n\leqslant p-$ 1 and $S=S_{n}(q)$ for $q=p^{m}$.
@@
    For $0\leqslant i\leqslant n$, $[C_{i},S]=C_{i-1}$. In particular, for $i\leqslant n-1$, $C_{i}$ is the $(n-i)$th term of the lower central series of $S$.
@@
Let $12\leqslant n\leqslant p-$ 1 and $q=p^{m}$. The following hold:
@@
$=$ $((1,1,\left(\begin{smallmatrix}1&0\\ \lambda+\mu&1\end{smallmatrix}\right)),\displaystyle\sum_{i=0}^{n}\lambda_{i}x^{n-i}\left(\sum_{j=0}^{i}{i\choose j}(\mu x)^{j}y^{i-j}\right)+\mu_{i}x^{n-i}y^{i})$
@@
Let $S=S_{n}(q)$. By Lemma 2.4 (4), $V=V_{n}(q)$ is a characteristic subgroup of $S$ and therefore so is $[V,S,S]$. We observe by (2) that $S/[V,S,S]\cong S_{1}(q)$, which itself is isomorphic to a Sylow $p$-subgroup of $\operatorname{SL}_{3}(q)$ by (1). Notice that $C_{\mathrm{Aut}(S)}(S/[V,S,S])\leq C_{\mathrm{Aut}(S)}(S/\Phi(S))$ is a normal $p$-subgroup of $\mathrm{Aut}(S)$ and that $\mathrm{Aut}(S)/C_{\mathrm{Aut}(S)}(S/[V,S,S])$ acts faithfully on $S/[V,S,S]$. Since $\mathrm{Aut}(S)$ leaves $V$ invariant by Lemma 2.4 (4), we see that the action of $\mathrm{Aut}(S)$ on $S/[V,S,S]$ preserves an elementary abelian subgroup of order $q^{2}$ in $S/[V,S,S]$. Hence $\mathrm{Aut}(S)$ is solvable by Proposition 2.5. Moreover, the Hall $p^{\prime}$-subgroups of $\mathrm{Aut}(S)$ have order dividing $(q-1)^{2}.m_{p^{\prime}}$, where $m_{p^{\prime}}$ denotes the $p^{\prime}$-part of $m$. Finally, consider $P^{*}=P_{n}^{*}(q)$. We have $\mathrm{Aut}_{P^{*}}(S)=\mathrm{Aut}_{BV}(S)\cong BV/Z(S)C_{D^{*}}(V)$ which has order $q^{n+1}(q-1)^{2}.m_{p^{\prime}}$ by Lemma 2.2. Thus the upper bound of $(q-1)^{2}.m_{p^{\prime}}$ for the order of a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}(S)$ is attained by a subgroup of $\mathrm{Aut}_{P^{*}}(S)$. This proves (4). The final statement follows from Lemma 2.2 and the fact that $|N_{D^{*}}(S_{n}(q))/S_{n}(q)|=(q-1)^{3}.m_{p^{\prime}}$. ∎
@@
Suppose that $12\leqslant n\leqslant p-$ 1 and set $S=S_{n}(q)$, $V=V_{n}(q)$ and $L=\mathrm{Aut}_{P_{n}^{*}(q)}(S_{n}(q))$. Let $L_{0}=C_{L}(V/[V,S])$. Then $L_{0}$ acts and irreducibly on $C_{V}(S)$ considered as a $\mathbb{F}_{p}L_{0}$-module. Furthermore, the action is trivial if and only if $n=p-1=q-1$.
@@
$L_{0}=C_{L}(V/[V,S])=\left\{c_{x}\mid x=\left(1,1,\begin{pmatrix}ad^{-1}&0\\ c&1\end{pmatrix}\right)\in C_{D^{*}}(V/[V,S])\right\}.$
By Lemma 2.4(2), $C_{V}(S)=\langle x^{n}\rangle_{\mathbb{K}}$ and, writing $Z=C_{V}(S)$, we see $c_{x}:Z\rightarrow Z$ is multiplication by $\lambda^{n}$ for $\lambda\in\mathbb{K}^{*}$. Hence $L_{0}$ induces a group of order $(q-1)/\gcd(q-1,n)$ on $Z$. It follows that $Z$ is irreducible as a $\mathbb{F}_{p}L_{0}$-module and is the trivial module if and only if $n=p-1=q-1$. ∎
@@
Suppose that $S=S_{n}(q)$ with $q=p^{m}$ and $11\leqslant n\leqslant p-$ 1, and $V=V_{n}(q)\leq S$. Then, for $22\leqslant i\leqslant n-$ 2, we have $Z_{n-i+1}(S)=[V,S;i]\cong V_{n-i}(q)$ as an $\mathbb{F}_{p}U$-module. In particular, $[V,S;i]U\cong S_{n-i}(q)$.
@@
$[V,S]=[V,z]\cong V/C_{V}(z)=V/C_{V}(S)\cong V_{n-1}(q)$
as $\mathbb{F}_{p}U$-modules. It follows that $[V,S]U\cong S_{n-1}(q)$. Now iterate this argument to obtain the result. ∎
@@
Let $S=S_{n}(q)$ for $12\leqslant n\leqslant p-$ 1. Then $UZ_{2}(S)\cong S_{1}(q)$.
@@
Let $S=S_{n}(q)$ and $V=V_{n}(q)$ with $11\leqslant n\leqslant p-$ 1. Put $Z_{i}=Z_{i}(S)$ for $1\leqslant i\leqslant n$ and $Z_{n+1}=V$. Then for $1\leqslant i\leqslant n$ for $x\in Z_{i+1}\setminus Z_{i}$, we have $[x,S]=Z_{i}$.
@@
Suppose first that $n=1$. Then $x\in V\setminus Z_{1}$ and the result follows from an elementary calculation. Now suppose that $n>1$. Set $T=Z_{i+1}U$. Then $T\cong S_{i}(q)$ by Lemma 2.8. If $i+1<n$, the result then follows by induction. Thus we may assume that $T=S$ and $x\in V\setminus Z_{n}$. By induction, and using Lemma 2.6 (3), $Z_{n}=[x,S]Z_{1}$. Notice that $Z_{2}\leq[x,S]Z_{1}$ and so there exists $y\in([x,S]\cap Z_{2})\setminus Z_{1}$. Since $[x,S]$ is normalised by $U$, $[x,S]\geq[y,U]=[y,Z_{2}U]=Z_{1}$ by Lemma 2.8 and induction as $Z_{2}U\cong S_{1}(q)$. Hence $Z_{n}=[x,S]Z_{1}=[x,S]$ and this proves the claim. ∎
@@
    $[V,z]=[V,S]=\langle x^{i}y^{p-i}\mid 2\leqslant i\leqslant p\rangle_{\mathbb{K}}$, and in particular, $|V/[V,S]|=q^{2}$;
@@
from which we see that $\lambda^{p-1}=-\gamma^{-1}$ which only has $p-1$ solutions in $\mathbb{K}$. Hence $\dim_{\mathbb{K}}C_{V}(S)\neq 1$ if and only if $\mathbb{K}$ has $p$ elements. Applying a similar reasoning to before, when $q=p$ we witness that $x^{p-1}y-y^{p}\in C_{V}(S)$. This proves (1).
It is clear from the action of $S$ that $\langle x^{p},y^{p},x^{p-1}y\rangle_{\mathbb{K}}\leq Z_{2}(S)$. Moreover, from the structure of $V/W$ we must have that $Z_{2}(S)/W\leq\langle x^{p-1}y,x^{p-2}y^{2}\rangle_{\mathbb{K}}W/W$. Since $Z_{2}(S)$ is a $\mathbb{K}$-space, to show that $Z_{2}(S)=\langle x^{p},y^{p},x^{p-1}y\rangle_{\mathbb{K}}$ we need only verify that $x^{p-2}y^{2}\not\in Z_{2}(S)$. As before, recognise $g_{\lambda}$ with $\left(1,1\begin{pmatrix}1&0\\ \lambda&1\end{pmatrix}\right)$, and then $[x^{p-2}y^{2},g_{\lambda}]=\lambda^{2}x^{p}+2\lambda x^{p-1}y$ which is not contained in $C_{V}(S)$ for any $q$ by (1). Hence, $Z_{2}(S)$ is as described. Since $Z_{2}(S)/W=C_{V/W}(S)$, the observation that $Z_{p}(S)=S$ and $|Z_{i}(S)/Z_{i-1}(S)|=q$ whenever $2<i<p$ follow from the structure of $V/W$ as witnessed in Lemma 2.4. This proves (2).
For $z\in S\setminus V$, commutation by $z$ induces a homomorphism from $V$ to $[V,z]$ with kernel $C_{V}(z)$. By (1), we see that $|C_{V}(z)|=|V/[V,z]|\leqslant q^{2}$. We claim that $[V,S]=[V,z]$ has index $q^{2}$ in $V$. Observe that $[V,S]$ is generated by $\{[x^{i}y^{j},g_{\lambda}]\mid\lambda\in\mathbb{K}\}$. We will show that $[V,S]=\langle x^{i}y^{p-i}\mid i\geq 2\rangle_{\mathbb{K}}$, which has dimension $p-1$. First note that $[y^{p},g_{\lambda}]=\lambda^{p}x^{p}\in[V,S]$. Then, for $1\leqslant k\leqslant p$,
$[x^{k}y^{p-k},g_{\lambda}]$ $=$ $x^{k}(\lambda x+y)^{p-k}-x^{k}y^{p-k}=\sum_{j=1}^{p-k}{{p-k}\choose j}\lambda^{j}x^{j+k}y^{p-k-j}\in[V,S]$
@@
We have that $Z_{2}(S)/C_{V}(z)=WC_{V}(z)/C_{V}(z)\leq C_{V/C_{V}(z)}(z)$. As in the proof of (2), using the structure of $V/W$, to prove that $WC_{V}(z)/C_{V}(z)=C_{V/C_{V}(z)}(z)$ it suffices to show that $[x^{p-2}y^{2},z]\not\leq C_{V}(z)$. This is the same calculation as in the proof of (2), which yields the result. Since $W\leq Z_{2}(S)$, the remainder of (5) follows upon applying the results in Lemma 2.4 to $V/W$.
@@
Let $S=S_{\Lambda}(q)$ and $V=\Lambda(q)$. Set $W$ to be the $(p-1)$-dimensional $\mathbb{K}D$-submodule of $V$, when $V$ is regarded as a $\mathbb{K}D$-module, and let $z\in S\setminus V$. Then
@@
    $Z_{2}(S)$ has order $q^{3}$, $Z_{p}(S)=S$ and for $12\leqslant i\leqslant p-$ 1 we have that $|Z_{i}(S)/Z_{i-1}(S)|=q$;
@@
    if $q>p$ then we have that $|V/[V,S]|=q$, $|V/[V,z]|=q^{2}$, $[V,z]\cap W=[W,z]$ and $[V,S]=WC_{V}(S)=[V,z]C_{V}(S)=[V,z]W$;
@@
Let $S=S_{\Lambda}(q)$ where $q=p^{m}$ and $m>1$. Then $V:=\Lambda(q)$ and $U[V,S]$ are the unique normal subgroups of $S$ of index $q$ which have exponent $p$. In particular, $U[V,S]$ is characteristic in $S$.
@@
Let $a\in A\setminus(A\cap V)$. By [Hup67, Theorem 9.7], for all $v\in A\cap V$ and $a\in A$ we have that $[v,a;p-1]\in[V,S;p]=1$. By Lemma 2.12, we have that $[[V,S],a;p-1]=1$ so that $[[V,S](A\cap V),a;p-1]=1$ and again considering Lemma 2.12, we see that $A\cap V\leq[V,S]$. Since $|S/A|=q$ we must have that $[V,S]=A\cap V$, $A$ has nilpotency class $p-1$ and $S=AV$. Another application of [Hup67, Theorem 9.7] reveals that $S$ has exponent $p^{2}$.
Now, $A/W$ is elementary abelian of order $q^{2}$, as are both $U[V,S]/W$ and $V/W$. Since $S/W\cong S_{1}(q)$, $S/W$ has $q+1$ elementary abelian subgroups of order $q^{2}$ which contain $[V,S]/W$, and the preimages of these subgroups cover all elements of $S$. There is an element of order $q-1$ in $P_{\Lambda}^{*}(q)$ which normalises $U[V,S]$, $V$, $[V,S]$ and $W$ and permutes the remaining $q-1$ elementary abelian subgroups of $S/W$. Since $S$ has exponent $p^{2}$, the preimages in $S$ of the permuted subgroups all contain elements of order $p^{2}$, and finally, since $A\not\leq V$, we conclude that $A=U[V,S]$. That is, $U[V,S]$ is the unique normal subgroup of $S$ of exponent $p$ and index $q$ which is not equal to $V$. Since $V$ is characteristic in $S$, so too is $U[V,S]$. ∎
@@
$C_{L}(C_{[V,S,S]}(S))=\left\{c_{x}\mid x=\left(1,a^{-1}d,\begin{pmatrix}ad^{-1}&0\\ c&1\end{pmatrix}\right)\in C_{D^{*}}(C_{[V,S,S]}(S))\right\}.$
@@
$\left\{c_{x}\mid x=\left(1,a^{-1}d,\begin{pmatrix}ad^{-1}&0\\ 0&1\end{pmatrix}\right)\in C_{D^{*}}(C_{[V,S,S]}(S))\right\}.$
@@
for $k\in K_{1}$. Note that $ad^{-1}$ has multiplicative order at least $\frac{q-1}{2}$ in $\mathbb{K}$. Choose $k$ such that $K_{1}=\langle k\rangle$. By linearity $v\cdot k=v$ implies that $\lambda_{i}=0$ unless $i=p-1$ (since $q>p$, $(ad^{-1})^{i-p-1}=1$ only if $i=p-1$). Hence, $C_{V}(K_{1})=\langle\overline{xy^{p-1}}\rangle_{\mathbb{K}}$, as desired.
@@
$\left\{c_{x}\mid x=\left(1,\theta,\begin{pmatrix}a&0\\ 0&d\end{pmatrix}\right)\in D^{*}\right\}$
@@
For $S\in\{S_{n}(q),S_{\Lambda}(q)\}$ where $11\leqslant n\leqslant p-$ 1, we set $d=1$ if $S=S_{n}(q)$ and $d=q$ if $S=S_{\Lambda}(q)$. Define
$\begin{array}[]{rcl}\mathcal{B}(S)&=&\{B\leq S\mid\mbox{$B$ is elementary abelian, }|B|=q^{2}d\mbox{ and }S=BV\}\mbox{; and}\\ \mathcal{C}(S)&=&\{C\leq S\mid\mbox{$C$ is class 2 and exponent $p$, }|C|=q^{3}d\mbox{ and }S=CV\}.\end{array}$
Observe that $Z(S)\leq B$ for $B\in\mathcal{B}(S)$ and $Z_{2}(S)\leq C$ for $C\in\mathcal{C}(S)$. Note also that $\mathcal{C}(S)$ only makes sense for $S=S_{n}(q)$ when $n\geqslant 2$.

### v2

@@ 2. The structure of polynomial $p$-groups and their automorphism group
Observe that $C_{X}(C_{V_{n}(q)}(T))\leq N_{X}(T)$ and $|N_{X}(T)|=q.(q-1)$. Clearly, $T$ centralises $C_{V_{n}(q)}(T)$. Let $K$ be a Hall $p^{\prime}$-subgroup of $N_{X}(T)$ so that $K$ is cyclic of order q1q-1 and let $k$ be a generator for $K$. Then we can arrange that $k$ acts as $\lambda^{n}$ on $C_{V_{n}(q)}(T)$ for some $\lambda\in\mathbb{K}$ of multiplicative order q1q-1. Thus every element of $K$ acts as $\mu^{n}$ on $C_{V_{n}(q)}(T)$ for some distinct $\mu\in\mathbb{K}^{*}$. Hence, the elements of $K$ which centralise $C_{V_{n}(q)}(T)$ are in bijection with the $n$th roots of unity in $\mathbb{K}$, which have order (n,q1)(n,q-1) and the claim holds. ∎
For $n\leqslant p-1$ and $0\neq f=\displaystyle\sum_{i=0}^{n}a_{i}x^{n-i}y^{i}\in V$ define $\mathrm{wt}(f)=\mathrm{max}\{i\mid a_{i}\neq 0\}$ to be the weight of $f$. Also define $\mathrm{wt}(0)=-1$. For $1\leqslant i\leqslant n$, set $C_{i}=\{f\in V\mid\mathrm{wt}(f)\leqslant i\}.$ For each $1\leqslant i\leqslant n$, $C_{i}$ is a subgroup of $V_{n}(q)$ of order $q^{i+1}$. In particular, $C_{-1}=1$ and $C_{n}=V$.
@@
Suppose that $1\leqslant n\leqslant p-1$ and $S=S_{n}(q)$ for $q=p^{m}$.
@@
    For $0\leqslant i\leqslant n$, $C_{i},S]=C_{i-1}$. In particular, for $i\leqslant n-1$, $C_{i}$ is the (ni)(n-i)th term of the lower central series of $S$.
@@
Let $2\leqslant n\leqslant p-1$ and $q=p^{m}$. The following hold:
@@
$=$ $((1,1,\left(\begin{smallmatrix}1&0\\ \lambda+\mu&1\end{smallmatrix}\right)),\displaystyle\sum_{i=0}^{n}\lambda_{i}x^{n-i}\left(\sum_{j=0}^{i}\binom{i}{j}(\mu x)^{j}y^{i-j}\right)+\mu_{i}x^{n-i}y^{i})$
@@
Let $S=S_{n}(q)$. By Lemma 2.8 (4), $V=V_{n}(q)$ is a characteristic subgroup of $S$ and therefore so is $[V,S,S]$. We observe by (2) that $S/[V,S,S]\cong S_{1}(q)$, which itself is isomorphic to a Sylow $p$-subgroup of $\operatorname{SL}_{3}(q)$ by (1). Notice that $C_{\mathrm{Aut}(S)}(S/[V,S,S])\leq C_{\mathrm{Aut}(S)}(S/\Phi(S))$ is a normal $p$-subgroup of $\mathrm{Aut}(S)$ and that $\mathrm{Aut}(S)/C_{\mathrm{Aut}(S)}(S/[V,S,S])$ acts faithfully on $S/[V,S,S]$. Since $\mathrm{Aut}(S)$ leaves $V$ invariant by Lemma 2.8 (4), we see that the action of $\mathrm{Aut}(S)$ on $S/[V,S,S]$ preserves an elementary abelian subgroup of order $q^{2}$ in $S/[V,S,S]$. Hence $\mathrm{Aut}(S)$ is solvable by Proposition 2.9. Moreover, the Hall $p^{\prime}$-subgroups of $\mathrm{Aut}(S)$ have order dividing $(q-1)^{2}.m_{p^{\prime}}$, where $m_{p^{\prime}}$ denotes the $p^{\prime}$-part of $m$.
We have $\mathrm{Aut}_{P_{n}^{*}(q)}(S)=\mathrm{Aut}_{BV}(S)\cong BV/Z(S)C_{D^{*}}(V)$ which has order $q^{n+1}(q-1)^{2}.m_{p^{\prime}}$ by Lemma 2.6. Thus the upper bound of $(q-1)^{2}.m_{p^{\prime}}$ for the order of a Hall $p^{\prime}$-subgroup of $\mathrm{Aut}(S)$ is attained by a subgroup of $\mathrm{Aut}_{P_{n}^{*}(q)}(S)$. This proves (4). The final statement follows from Lemma 2.6 and the fact that $|N_{D^{*}}(S_{n}(q))/S_{n}(q)|=(q-1)^{3}.m_{p^{\prime}}$. ∎
@@
Suppose that $2\leqslant n\leqslant p-1$ and set $S=S_{n}(q)$, $V=V_{n}(q)$ and $L=\mathrm{Aut}_{P_{n}^{*}(q)}(S_{n}(q))$. Let $L_{0}=C_{L}(V/[V,S])$. Then $L_{0}$ acts and irreducibly on $C_{V}(S)$ considered as a $\mathbb{F}_{p}L_{0}$-module. Furthermore, the action is trivial if and only if n=p1=q1n=p-1=q-1.
@@
$L_{0}=C_{L}(V/[V,S])=\left\{c_{z}\mid z=\left(1,1,\begin{pmatrix}ad^{-1}&0\\ c&1\end{pmatrix}\right)\in C_{D^{*}}(V/[V,S])\right\}.$
By Lemma 2.8(2), $C_{V}(S)=\langle x^{n}\rangle_{\mathbb{K}}$ and, writing $Z=C_{V}(S)$, we see $c_{z}:Z\rightarrow Z$ is multiplication by $\lambda^{n}$ for $\lambda\in\mathbb{K}^{*}$. Hence $L_{0}$ induces a group of order $(q-1)/\gcd(q-1,n)$ on $Z$. It follows that $Z$ is irreducible as a $\mathbb{F}_{p}L_{0}$-module and is the trivial module if and only if n=p1=q1n=p-1=q-1. ∎
@@
Suppose that $S=S_{n}(q)$ with $q=p^{m}$ and $1\leqslant n\leqslant p-1$, and $V=V_{n}(q)\leq S$. Then, for $2\leqslant i\leqslant n-2$, we have $Z_{n-i+1}(S)=[V,S;i]\cong V_{n-i}(q)$ as an $\mathbb{F}_{p}U$-module. In particular, $V,S;i]U\cong S_{n-i}(q)$.
@@
$V,S]=[V,z]\cong V/C_{V}(z)=V/C_{V}(S)\cong V_{n-1}(q)$
as $\mathbb{F}_{p}U$-modules. It follows that $V,S]U\cong S_{n-1}(q)$. Now iterate this argument to obtain the result. ∎
@@
Let $S=S_{n}(q)$ for $2\leqslant n\leqslant p-1$. Then $UZ_{2}(S)\cong S_{1}(q)$.
@@
Let $S=S_{n}(q)$ and $V=V_{n}(q)$ with $1\leqslant n\leqslant p-1$. Put $Z_{i}=Z_{i}(S)$ for $1\leqslant i\leqslant n$ and $Z_{n+1}=V$. Then for $1\leqslant i\leqslant n$ for $z\in Z_{i+1}\setminus Z_{i}$, we have $z,S]=Z_{i}$.
@@
Suppose first that $n=1$. Then $z\in V\setminus Z_{1}$ and the result follows from an elementary calculation. Now suppose that $n>1$. Set $T=Z_{i+1}U$. Then $T\cong S_{i}(q)$ by Lemma 2.12. If $i+1<n$, the result then follows by induction. Thus we may assume that $T=S$ and $z\in V\setminus Z_{n}$. By induction, and using Lemma 2.10 (3), $Z_{n}=[z,S]Z_{1}$. Notice that $Z_{2}\leq[z,S]Z_{1}$ and so there exists $y\in([z,S]\cap Z_{2})\setminus Z_{1}$. Since $[z,S]$ is normalised by $U$, $z,S]\geq[y,U]=[y,Z_{2}U]=Z_{1}$ by Lemma 2.12 and induction as $Z_{2}U\cong S_{1}(q)$. Hence $Z_{n}=[z,S]Z_{1}=[z,S]$ and this proves the claim. ∎
@@
    $V,z]=[V,S]=\langle x^{i}y^{p-i}\mid 2\leqslant i\leqslant p\rangle_{\mathbb{K}}$, and in particular, $|V/[V,S]|=q^{2}$;
@@
from which we see that $\lambda^{p-1}=-\gamma^{-1}$ which only has p1p-1 solutions in $\mathbb{K}$. Hence $\dim_{\mathbb{K}}C_{V}(S)\neq 1$ if and only if $\mathbb{K}$ has $p$ elements. Applying a similar reasoning to before, when $q=p$ we witness that $x^{p-1}y-y^{p}\in C_{V}(S)$. This proves (1).
It is clear from the action of $S$ that $\langle x^{p},y^{p},x^{p-1}y\rangle_{\mathbb{K}}\leq Z_{2}(S)$. Moreover, from the structure of $V/W$ we must have that $Z_{2}(S)/W\leq\langle x^{p-1}y,x^{p-2}y^{2}\rangle_{\mathbb{K}}W/W$. Since $Z_{2}(S)$ is a $\mathbb{K}$-space, to show that $Z_{2}(S)=\langle x^{p},y^{p},x^{p-1}y\rangle_{\mathbb{K}}$ we need only verify that $x^{p-2}y^{2}\not\in Z_{2}(S)$. As before, recognise $g_{\lambda}$ with $\left(1,1\begin{pmatrix}1&0\\ \lambda&1\end{pmatrix}\right)$, and then $x^{p-2}y^{2},g_{\lambda}]=\lambda^{2}x^{p}+2\lambda x^{p-1}y$ which is not contained in $C_{V}(S)$ for any $q$ by (1). Hence, $Z_{2}(S)$ is as described. Since $Z_{2}(S)/W=C_{V/W}(S)$, the observation that $Z_{p}(S)=S$ and $|Z_{i}(S)/Z_{i-1}(S)|=q$ whenever $2<i<p$ follow from the structure of $V/W$ as witnessed in Lemma 2.8. This proves (2).
For $z\in S\setminus V$, commutation by $z$ induces a homomorphism from $V$ to $[V,z]$ with kernel $C_{V}(z)$. By (1), we see that $|C_{V}(z)|=|V/[V,z]|\leqslant q^{2}$. We claim that $[V,S]=[V,z]$ has index $q^{2}$ in $V$. Observe that $[V,S]$ is generated by $\{[x^{i}y^{j},g_{\lambda}]\mid\lambda\in\mathbb{K}\}$. We will show that $V,S]=\langle x^{i}y^{p-i}\mid i\geq 2\rangle_{\mathbb{K}}$, which has dimension p1p-1. First note that $y^{p},g_{\lambda}]=\lambda^{p}x^{p}\in[V,S]$. Then, for $1\leqslant k\leqslant p$,
$[x^{k}y^{p-k},g_{\lambda}]$ $=$ $x^{k}(\lambda x+y)^{p-k}-x^{k}y^{p-k}=\sum_{j=1}^{p-k}\binom{p-k}{j}\lambda^{j}x^{j+k}y^{p-k-j}\in[V,S]$
@@
We have that $Z_{2}(S)/C_{V}(z)=WC_{V}(z)/C_{V}(z)\leq C_{V/C_{V}(z)}(z)$. As in the proof of (2), using the structure of $V/W$, to prove that $WC_{V}(z)/C_{V}(z)=C_{V/C_{V}(z)}(z)$ it suffices to show that $x^{p-2}y^{2},z]\not\leq C_{V}(z)$. This is the same calculation as in the proof of (2), which yields the result. Since $W\leq Z_{2}(S)$, the remainder of (5) follows upon applying the results in Lemma 2.8 to $V/W$.
@@
Let $S=S_{\Lambda}(q)$ and $V=\Lambda(q)$. Set $W$ to be the (p1)(p-1)-dimensional $\mathbb{K}D$-submodule of $V$, when $V$ is regarded as a $\mathbb{K}D$-module, and let $z\in S\setminus V$. Then
@@
    $Z_{2}(S)$ has order $q^{3}$, $Z_{p}(S)=S$ and for $2\leqslant i\leqslant p-1$ we have that $|Z_{i}(S)/Z_{i-1}(S)|=q$;
@@
    if $q>p$ then we have that $|V/[V,S]|=q$, $|V/[V,z]|=q^{2}$, $V,z]\cap W=[W,z]$ and $V,S]=WC_{V}(S)=[V,z]C_{V}(S)=[V,z]W$;
@@
Let $S=S_{\Lambda}(q)$ where $q=p^{m}$ and $m>1$. Then $V=\Lambda(q)$ and $U[V,S]$ are the unique normal subgroups of $S$ of index $q$ which have exponent $p$. In particular, $U[V,S]$ is characteristic in $S$.
@@
Let $a\in A\setminus(A\cap V)$. By [HUP67, Theorem 9.7], for all $v\in A\cap V$ and $a\in A$ we have that $v,a;p-1]\in[V,S;p]=1$. By Lemma 2.16, we have that [[V,S],a;p1]=1[[V,S],a;p-1]=1 so that $V,S](A\cap V),a;p-1]=1$ and again considering Lemma 2.16, we see that $A\cap V\leq[V,S]$. Since $|S/A|=q$ we must have that $V,S]=A\cap V$, $A$ has nilpotency class p1p-1 and $S=AV$. Another application of [HUP67, Theorem 9.7] reveals that $S$ has exponent $p^{2}$.
Now, $A/W$ is elementary abelian of order $q^{2}$, as are both $U[V,S]/W$ and $V/W$. Since $\cong S_{1}(q), S/WS/W$ has $q+1$ elementary abelian subgroups of order $q^{2}$ which contain $[V,S]/W$, and the preimages of these subgroups cover all elements of $S$. There is an element of order q1q-1 in $P_{\Lambda}^{*}(q)$ which normalises $U[V,S]$, $V$, $[V,S]$ and $W$ and permutes the remaining q1q-1 elementary abelian subgroups of $S/W$. Since $S$ has exponent $p^{2}$, the preimages in $S$ of the permuted subgroups all contain elements of order $p^{2}$, and finally, since $A\not\leq V$, we conclude that $A=U[V,S]$. That is, $U[V,S]$ is the unique normal subgroup of $S$ of exponent $p$ and index $q$ which is not equal to $V$. Since $V$ is characteristic in $S$, so too is $U[V,S]$. ∎
@@
$C_{L}(C_{[V,S,S]}(S))=\left\{c_{z}\mid z=\left(1,a^{-1}d,\begin{pmatrix}ad^{-1}&0\\ c&1\end{pmatrix}\right)\in C_{D^{*}}(C_{[V,S,S]}(S))\right\}.$
@@
$\left\{c_{z}\mid z=\left(1,a^{-1}d,\begin{pmatrix}ad^{-1}&0\\ 0&1\end{pmatrix}\right)\in C_{D^{*}}(C_{[V,S,S]}(S))\right\}.$
@@
for $k\in K_{1}$. Note that $ad^{-1}$ has multiplicative order at least $\frac{q-1}{2}$ in $\mathbb{K}$. Choose $k$ such that $K_{1}=\langle k\rangle$. By linearity $v\cdot k=v$ implies that $\lambda_{i}=0$ unless i=p1i=p-1 (since $q>p$, $(ad^{-1})^{i-p-1}=1$ only if i=p1i=p-1). Hence, $C_{V}(K_{1})=\langle\overline{xy^{p-1}}\rangle_{\mathbb{K}}$, as desired.
@@
$\left\{c_{z}\mid z=\left(1,\theta,\begin{pmatrix}a&0\\ 0&d\end{pmatrix}\right)\in D^{*}\right\}$
@@
Notation 2.21.
For $S\in\{S_{n}(q),S_{\Lambda}(q)\}$ where $1\leqslant n\leqslant p-1$, we set $d=1$ if $S=S_{n}(q)$ and $d=q$ if $S=S_{\Lambda}(q)$. Set $R:=UZ(S)$ and $Q:=UZ_{2}(S)$.
Define
[…]
Then $R\in\mathcal{B}(S)$ and $Q\in\mathcal{C}(S)$.
Observe that $Z(S)\leq B_{0}$ for $B_{0}\in\mathcal{B}(S)$ and $Z_{2}(S)\leq C_{0}$ for $C_{0}\in\mathcal{C}(S)$. Note also that $\mathcal{C}(S)$ only makes sense for $S=S_{n}(q)$ when $n\geqslant 2$.